A chemistry student wishes to make 100 c.c of 27% acid solution. he has two solutions of 20% acid and 30% acid. how many cubic centimeters of each solution should be combined to get the required solution?
1
Expert's answer
2018-03-30T08:09:44-0400
x - c.c. 20% acid
y - c.c 30% acid

x + y = 100
(0.2x + 0.3y)/100 = 0.27

System of linear equations

x + y = 100
x + 1.5y = 135

x= 30 y = 70